OpenStudy (anonymous):

Solve 10 - 12x^2 = 1 by inspection.

OpenStudy (kamibug):

\(\Huge\cal\color{seagreen}{Welcome~to~Openstudy!}\)

OpenStudy (kamibug):

To solve for x you need to isolate it on one side of the equal sign. To do that you can use inverse variations. 10-12x^2 = 1 -12x^2 = -9 x^2 = 9/12 9/12 = 3/4 x = the square root of 3/4 or 0.75 Hope this helps! :)

OpenStudy (anonymous):

thank you!!